Jack needs to prepare 12 ounces of a 9% hydrochloric acid solution. the only solution she currently has avaliable are 4% and 12% hydrochloric acid. How much should she use of each solution to obtain the needed solution?
Answer by checkley77(12844)

.12x+.04(12-x)=.09*12
.12x+.48-.04x=1.08
.08x=1.08-.48
.08x=.6
x=.6/.08
x=7.5 oz. of 12% acid is used.
12-7.5=4.5 oz of 4% acid is used.
proof:
.12*7.5+.04*4.5=1.08
.9+.18=1.08
1.09=1.08